OpenStudy (anonymous):

help. medal and fan c: Simplify 14+1/x2 -------- 28x+ 2/x

OpenStudy (anonymous):

a. 5/2x b. 1/2x c. x/2 d. 2x

OpenStudy (anonymous):

\[\frac{ 14+\frac{ 1 }{ x ^{2} } }{28x+\frac{ 2 }{ x} }=\frac{ \frac{ 14x ^{2}+1 }{ x ^{2} } }{\frac{ 28x ^{2}+2 }{x} }\] \[=\frac{ 14x ^{2}+1 }{ x ^{2} }*\frac{ x }{2\left( 14x ^{2}+1 \right) }=\frac{ 1 }{ 2x }\]

OpenStudy (anonymous):

thanks man. c: